OpenStudy (anonymous):

Find the rectangular coordinates for the point whose polar coordinates are given. (√3, -5π/3)

OpenStudy (anonymous):

lols

OpenStudy (anonymous):

so -5pi/3 is equal to pi/3

OpenStudy (anonymous):

but better check with someone else before u enter

OpenStudy (anonymous):

so x= sqrt(3) cos(pi/3) , y= sqrt(3) sin(pi/3) x= (1/2)sqrt(3) , y = (3/2)

OpenStudy (anonymous):

but dont enter yet lols

OpenStudy (anonymous):

haha.. ok

OpenStudy (anonymous):

u tell me when to

OpenStudy (anonymous):

it looks correct, you want check it yourself if you want, bit of practice in going the other way

OpenStudy (anonymous):

plot the point, and then try to convert it to polar form , if you get what you started with then its right ( assuming no silly algebra mistakes )

OpenStudy (anonymous):

yeah, it's right :O